Circular No. 6782

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE N. Regnault and O. Perdereau, Orsay, on behalf of the EROS collab-

oration (cf. IAUC 6605, 6628, 6760), report the discovery of four apparent supernovae in their systematic search with a 1^{o}-square CCD camera on the Marly 1-m telescope at the European Southern Observatory (ESO):

SN  1997 UT  α_{2000}  δ_{2000}  V  Offset
1997eb Nov. 19 2^{h}53^{m}00.^{s}15 − 3^{o}36′38′′.1 21  2′′.4 W, 1′′.8 N
1997ec Nov. 20 3 49 07.82 − 1 44 26.7 21  1′′.8 W, 2′′.7 N
1997ed Nov. 22 3 46 56.52 + 0 41 45.4 21  1′′.2 E, 2′′.4 N
1997ee Nov. 23 3 07 24.82 − 3 10 26.7 20.5 3′′.6 W, 7′′.8 S

All four supernovae were confirmed on EROS images obtained on Nov. 21, 23, or 24, and also on images taken with the ESO 1.5-m Danish telescope on Nov. 25 or 27. SNe 1997eb, 1997ec, and 1997ee were discovered in frames constructed from the differences between CCD frames taken on 1997 Oct. 7 and the discovery date; SN 1997ed appeared in a frame constructed from the difference between CCD frames taken on 1997 Oct. 24 and Nov. 22. The limiting magnitude of the search is ∼ 22, and the supernovae’s magnitudes above are approximate. Finding charts may be found on the EROS web

server at http://www.lal.in2p3.fr/EROS/SNe/Charts/.

COMET C/1997 O1 (TILBROOK)

Ephemeris from orbital elements on MPC 30632:

1997 /98 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 28 15^{h}05.^{m}25 +21^{o}19.^{′}2 2.907 2.332 45.^{o}9 17.^{o}7 14.0
Dec. 8 15 16.32 +24 16.5 2.892 2.433  53.2 18.9 14.2
18 15 27.08 +27 36.9 2.868 2.535  60.6 19.8 14.3
28 15 37.39 +31 22.4 2.840 2.638  68.2 20.2 14.5
Jan. 7 15 47.03 +35 33.8 2.815 2.740  75.5 20.3 14.6
17 15 55.72 +40 09.4 2.798 2.843  82.5 20.1 14.8
27 16 03.06 +45 05.7 2.795 2.945  88.9 19.5 14.9
Feb. 6 16 08.50 +50 16.2 2.811 3.047  94.2 18.8 15.1
16 16 11.24 +55 31.9 2.851 3.149  98.3 18.1 15.3
26 16 10.11 +60 42.1 2.915 3.250 100.8 17.4 15.4
Mar. 8 16 03.32 +65 35.1 3.005 3.351 101.7 16.9 15.6
1997 December 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green